the measure of an angle is 12 degrees less than twice the measure of its supplement. what is the measure of the angle?
Answer by richwmiller(17219)

a=s-12
Now the question- what is a supplementary angle? Do they add up to 90 or 180?
180=a+s
a=s-12
angle a 84 and supplement s 96 degrees
Reminder:
You are limited to one problem per submission.
No similar problems.
No more than 4 submissions daily